Calculate from the system
a) Ct
b) Qt when
c) Vab = 12 volts
d) Charge Q for each capacitor
3 Calcular de Sistema a. Ct b. Qr cuando c. Vn° 12 voltios 8. d. Carga Q por cada capacitor
Answer #1

a)= Ct=1/3\muf+1/6\muf= 2\muf

2+2=4\muf

b)= Qt=Ct x Vab

= 4\muf x 12 = 48 C

c) charge Q per capaciter

Q1=C1 X V = 3 X 12 = 36 C

Q2= C2 X V = 6 X 12 = 72 C

Q3= C3 X V = 2 X 12 =24 C